OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

@surjithayer please help

OpenStudy (anonymous):

@surjithayer please help please please please

OpenStudy (anonymous):

here slope is 4 eq through point (a,b) is y-b=4(x-a)

OpenStudy (anonymous):

so am I looking for the x or y axis points

OpenStudy (anonymous):

point is given (-3,-1)

OpenStudy (anonymous):

so fill in the blanks

OpenStudy (anonymous):

a=-3,b=-1

OpenStudy (anonymous):

y-1=4(x-3)

OpenStudy (anonymous):

y-(-1)=? for simplicity i have used a,b otherwise it is \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

-1y?

OpenStudy (anonymous):

no,y-(-1)=4(x-(-3)) y+1=4(x+3)

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thank you very much do you need help with any thing

OpenStudy (anonymous):

yw. where do you live?

OpenStudy (anonymous):

what does yw. mean and I live in pompano beach florida you?

OpenStudy (anonymous):

you are welcome.NJ

OpenStudy (anonymous):

your from New Jersey

OpenStudy (anonymous):

and y do u ask

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what does NJ mean then

OpenStudy (anonymous):

New Jersey